An air compressor has an inlet air enthalpy of 35 Btu/lb and an exit enthalpy of 70 Btu/lb. The mass flow of air is 3 lbs/s. If heat loss is 466.62 Btu/min, find the work input to the compressor.


1
Expert's answer
2021-09-29T17:50:51-0400

"H_1+Q=H_2+W,"

"W=Q+\Delta H,"

"W=466.62+3(70-35)60=6766.62~\frac{Btu}{min},"

"W=159.5~hp."